Exercise 7-8

On January 1, 2015, P Company acquired a 90% interest in S Company. During 2016, S Company sold merchandise to P Company at 25% above cost in the amount (selling price) of $241,800. At the end of the year, P Company had in its inventory one-third of the amount of goods purchased from S Company.

On January 1, 2016, P Company sold equipment that had a book value of $83,600 to S Company for $118,100. The equipment had an estimated remaining life of four years.

S Company reported net income of $116,100, and P Company reported net income of $293,000 from their independent operations (including sales to affiliates) for the year ended December 31, 2016.

Calculate controlling interest in consolidated net income for the year ended December 31, 2016.

How do you calculate the controlling interest in consolidated net income for Exercise 7-8?
Answer #1

% of holding interest in S Company by P company is 90%

Note 1

S Company sold merchandise to its holding company P company at 25% above cost

Selling price of inventory held by P co= 1/3*241800 = 80600

Less: cost of inventory held in P co = 1/3 of (241800/125%) = 64480

unrealised  profit = 16120

Note 2

profit reserved on sale of plant by P Co to S co = $118100-$83600

= $34500

Depreciation on profit reserved = $34500/4 = $8625

Note 3

Controlling interest in S ltd profit

Net income   = $116100

Less:Unrealised profit    = $16120

Net income for consolidation= $99980

P Co share @ 90%    = 89982

Calculation of consolidated profit of P co

Net income of P Co = $293000

Less Depreciation on profit reserved= $ 8625

Net income for consolidation = $ 284375

Add: Share of income from S co = $ 89982

consolidated net income = $374357